Two cars start at the same time, but travel in opposite directions. One car's average speed is 60 miles per hour (mph). At the end of 5 hours, the two cars are 450 miles apart. Find the average speed in mph of the other car.
Answer by ikleyn(52781)

.

Total distance equation


    5*60 + 5*x = 450 miles


    300 + 5x = 450

          5x = 450 - 300 = 150

           x             = 150/5 = 30  miles per hour.    ANSWER

Solved.